Circular No. 8029

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2002 X2 (NEAT) K. J. Lawrence reports the discovery of a comet by the NEAT team

(which now includes also R. Bambery, E. Helin, S. Pravdo, M. Hicks, and R. Thicksten) from CCD images taken with the 1.2-m Schmidt telescope at Palomar (discovery observation below). Following posting on the NEO Confirmation Page, J. Tichá and M. Tichý found the comet to show a 7′′ coma and a 14′′ tail in p.a. 45^{o} on Kleť images taken on Dec. 8.8 UT. P. Holvorcem that the co-addition of three 120-s exposures taken with the Tenagra II telescope at Nogales, AZ, on Dec. 8.20 shows a coma with diameter ∼ 7′′ and a tail ∼ 14′′ long in p.a. ∼ 55^{o}. F. B. Zoltowski, Edgewood, NM, notes that CCD images taken with a 0.3-m reflector on Dec. 9.15 shows the comet as diffuse with a 15′′ coma diameter and a possible faint tail in p.a. 240^{o}.

2002 UT  α_{2000}  δ_{2000}  m_{1}
Dec. 7.17100  2^{h}18^{m}17.^{s}91  −7^{o}46′37′′.0  18.4

Further astrometry (including prediscovery observations by LONEOS on Oct. 5 and Dec. 6, and by NEAT on Nov. 5) and the following orbital elements by B. G. Marsden appear on MPEC 2002-X51.

T = 2003 Mar. 29.5651 TT  ω = 356.^{o}0279
e = 0.373384  Ω = 78.1561 }2000.0
q = 2.528633 AU  i = 25.3531
a = 4.035377 AU n^{o} = 0.1215844 P = 8.106 years

XTE J1908+094 M. P. Rupen, A. J. Mioduszewski, and V. Dhawan, National Radio

Astronomy Observatory, report the detection of renewed radio emission associated with the x-ray transient and black-hole candidate XTE J1908+094 (IAUC 7856): “After the initial radio detection (IAUC 7874) on Mar. 21, observations at irregular intervals with the Very Large Array at 8.46 GHz gave detections through May 13, then non-detections between June 24 and Nov. 11 (to rms noise levels of 0.04–0.08 mJy/beam). On Dec. 7.99 UT, we again obtained a clear detection of 0.34 ± 0.055 mJy, at the position of the transient. Observations at other wavelengths, particularly x-ray and near-infrared (cf. IAUC 7897), are urgently requested.”

2002 December 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green